OpenStudy (anonymous):

What is the general form of the equation of a circle with a center at the (4, –2) and a radius of 9?

OpenStudy (anonymous):

A(x – 4)^2 + (y – 2)^2 = 81 B (x – 4)^2 + (y + 2^)2 = 81 C(x + 4^)2 + (y – 2)^2 = 81 D(x + 4)^2 + (y + 2)^2 = 81

OpenStudy (aravindg):

The general equation of a circle with centre (h,k) and radius r is \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (aravindg):

Use that here

OpenStudy (anonymous):

(x-4)^2+(y+2)^2=9^2

OpenStudy (anonymous):

(x^2+8)+(y^2+4)=81

OpenStudy (anonymous):

oh, 16?

OpenStudy (anonymous):

(x^2+16)+(y^2+4)=81

OpenStudy (aravindg):

(x-4)^2+(y+2)^2=9^2 IS RIGHT

OpenStudy (aravindg):

9^2=81

OpenStudy (anonymous):

ya that's what I did

OpenStudy (anonymous):

Now what do I do?

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

Nothing left !

OpenStudy (aravindg):

You got the answer :)

OpenStudy (anonymous):

OH, LOL thanks!

OpenStudy (aravindg):

lol Option B :)